OpenStudy (anonymous):

if x=5-2sqrt(6) then find the value x^2+1/x^2

OpenStudy (anonymous):

@amistre64 ....... can u pls.... check this out

OpenStudy (anonymous):

@ajprincess ....... can u pls.... check this out

OpenStudy (ajprincess):

put 5-2sqrt(6) in place of x and find the value

OpenStudy (anonymous):

im getting a little a troubled in between can u pls...do the problem with every step....pls....

OpenStudy (amistre64):

we would perfer not to "do all the work"; but to be able to guide you thru the process, interacting with youalong the way

OpenStudy (amistre64):

is this\[x^2+\frac1{x^2}\]or\[\frac{x^2+1}{x^2}\]

OpenStudy (anonymous):

im sry for the trouble but 2morrow is my exzam..and i dnt hv much time,,

OpenStudy (anonymous):

its ..\[x ^{2}+\frac{ 1 }{ x ^{2} }\]

OpenStudy (ajprincess):

\[(x+\frac{1}{x})^2\] \[=x^2+2*x*\frac{1}{x}+\frac{1}{x^2}\] \[=x^2+2+\frac{1}{x^2}\] \[(x+\frac{1}{x})^2-2=x^2+\frac{1}{x^2}\] Nw put 5-2sqrt(6) for x in \[(x+\frac{1}{x})^2-2 and try to find the value.

OpenStudy (ajprincess):

Nw put 5-2sqrt(6) for x in \[(x+\frac{1}{x})^2-2\] and try to find the value.

OpenStudy (amistre64):

and \(x=5-2\sqrt{6}\) so every place we see an "x" we plug in its value, right?

OpenStudy (amistre64):

i think aj might have misread the notation. which is common when trying to decipher the ascii text

OpenStudy (amistre64):

since this is the form you picked \[x^2+\frac1{x^2}\]lets make life simpler by working the general setup first; then we can insert the specifics \[x^2\frac{x^2}{x^2}+\frac1{x^2}\] \[\frac{x^4+1}{x^2}\] \[\frac{(5-2\sqrt{6})^4+1}{(5-2\sqrt{6})^2}\]

OpenStudy (amistre64):

your other problems gave you approximations for sqrts; does this one include approximations for sqrt(6) ??

hartnn (hartnn):

@amistre64 , i think aj's method works just fine and is simpler, can u plz check it again ??

OpenStudy (amistre64):

let me see :)

OpenStudy (amistre64):

aj apparently read it correctly, good job :)

OpenStudy (ajprincess):

Thanx a lot for the compliment:)

OpenStudy (ajprincess):

Can u find the value @basith?

OpenStudy (anonymous):

giv me a min....

OpenStudy (ajprincess):

ha k.

OpenStudy (anonymous):

ill do it here step by step ....pls see if it is correct...ok @ajprincess

OpenStudy (anonymous):

\[(5-2\sqrt{6}+\frac{ 1 }{ 5-\sqrt{6} })^{2}-2\]

OpenStudy (anonymous):

\[\frac{ 25-20\sqrt{6}+24 }{ 5+2\sqrt{6} }+\frac{ 1 }{ 25-20\sqrt{6}+24 }\]

OpenStudy (anonymous):

then wat????????? and it is nt 1 it is\[5+2\sqrt{6}\]

OpenStudy (ajprincess):

\[(5-2\sqrt6+\frac{ 1 }{ 5-2\sqrt6 })^2-2\] \[=(\frac{ 25-20\sqrt6+24 }{5-2\sqrt6}+\frac{ 1 }{ 5-2\sqrt6})^2-2\] \[=(\frac{ 25-20\sqrt6+24 + 1 }{ 5-2\sqrt6 })^2-2\]

OpenStudy (ajprincess):

Can u continue from this @basith?

OpenStudy (anonymous):

????......wat 2 do next..??? :(

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

Simplify the numerator and factorise it

OpenStudy (anonymous):

\[(\frac{ 50-20\sqrt{6} }{ 5-2\sqrt{6} })^{2}-2\]

OpenStudy (ajprincess):

yup right. Nw try to factorise the numerator.

OpenStudy (anonymous):

is that = \[(10)^{2}-2\] @ajprincess

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

so...98

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

i hv 1 mor question....pls wait

OpenStudy (anonymous):

if x+1/x=7 , then find the value of x^3+1/x^3

OpenStudy (anonymous):

i sthe answer just 343

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):

May I knw hw u got that answer?

OpenStudy (anonymous):

\[(x+\frac{ 1 }{ x })^{3}=(7)^{3}=343\]

OpenStudy (anonymous):

it is really wrong....isn't it.. :(

OpenStudy (ajprincess):

hmm ya. \[x^2+\frac{1}{x^2}=(x+\frac{1}{x})^2-2\] \[x^3+\frac{1}{x^3}=(x+\frac{1}{x})(x^2-1+\frac{1}{x^2})\] \[x^3+\frac{1}{x^3}=(x+\frac{1}{x})((x+\frac{1}{x})^2-2-1)\] \[=(x+\frac{1}{x})((x+\frac{1}{x})^2-2-1)\] \[=(x+\frac{1}{x})((x+\frac{1}{x})^2-3)\] Hope u can do it nw.

OpenStudy (anonymous):

wat did u di in the 3rd step???

OpenStudy (ajprincess):

can u plz type it down?

hartnn (hartnn):

@basith a simpler way to do that will be to use: \(\large (a+b)^3=a^3+b^3+3ab(a+b)\) here put a= x and b = 1/x so ab = x* (1/x) = 1 a+b = x+1/x = 7 can u solve further ??

OpenStudy (anonymous):

actually wat did u write first.....\[x ^{3}+\frac{ 1 }{ x ^{3} }=(x+\frac{ 1 }{ ^{3} })^{2}-2\] ????????????

OpenStudy (anonymous):

@ajprincess

OpenStudy (ajprincess):